OpenStudy (anonymous):

4(n+1)-4n=2(2-n)+2n? A.n=0 B.n=4 C.no solution D.infinite solution

OpenStudy (anonymous):

Are you trying to solve for 'n'?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Distribute what?

OpenStudy (anonymous):

example of distribution: 4 (n+1) = 4n + 4

OpenStudy (anonymous):

-3?

OpenStudy (anonymous):

4(n + 1) - 4n = 2(2 - n) + 2n (multiply through the parenthesis) 4n + 4 - 4n = 4 - 2n + 2n (combine like terms) 4 = 4....you can't solve for n so I am guessing there is no solution.

OpenStudy (anonymous):

n=0

OpenStudy (anonymous):

are you sure ...lets plug 0 in the equation and see.... 4(n + 1) - 4n = 2(2 - n) + 2n 4(0 + 1) - 4(0) = 2(2 - 0) + 2(0) 4(1) - 0 = 2(2) + 0 4 = 4 Your right....n = 0

OpenStudy (anonymous):

Thank you so much